# A Necessary Condition for the Submergence of Proto-Neutron Star Magnetic Fields by Supernova Fallback

### v1

@@ Abstract
Central compact objects (CCOs) are a subclass of neutron stars with a dipole magnetic field strength considerably weaker than those of radio pulsars and magnetars. One possible explanation for such weak magnetic fields in the CCOs is the hidden magnetic field scenario, in which supernova fallback submerges the magnetosphere of a proto-neutron star beneath a newly formed crust. However, the fallback mass and timescale required for this submergence process remain uncertain. We perform one-dimensional general relativistic magnetohydrodynamic simulations of the supernova fallback onto a magnetized proto-neutron star, while considering neutrino cooling. In our simulations, the infalling material compresses the magnetic field and drives a strong shock. The shock initially expands outward, but eventually stalls and recedes as neutrino cooling becomes significant. After the shock stalls, the gas density above the magnetosphere increases rapidly, potentially leading to the formation of a new crust. To understand the shock dynamics, we develop semi-analytic models that describe the resulting magnetospheric and shock radii when the shock stalls. By comparing the fallback time scale with the shock stalling time scale, corresponding to the waiting time for the new crust formation, we derive a necessary condition for the submergence of the PNS magnetic field. Our results will provide guidance for investigating the diversity of young isolated neutron stars through multidimensional simulations.
@@ 1 Introduction
Based on multi-wavelength observations, young isolated neutron stars (NSs) in our galaxy with ages of $\lesssim 1-10~{}{\rm kyr}$ are classified as magnetars, radio pulsars, and central compact objects (CCOs). The central engines powering these objects are believed to be the magnetic energy, rotation energy, and thermal energy of the NSs, respectively (e.g., see reviews by Enoto et al., 2019; Borghese & Esposito, 2023). Their characteristic values of the dipole magnetic field strength at the NS surface are considerably different: $>10^{14}~{}{\rm G}$ for magnetars, $10^{12-13}~{}{\rm G}$ for radio pulsars, and $<10^{11}~{}{\rm G}$ for CCOs. However, the origin of such diversity remains unclear.
What determines the magnetic field strength of NSs is an open question (e.g., see reviews by Igoshev et al., 2021). When a proto-NS (PNS) forms in a core-collapse supernova, the magnetic field of the progenitor core is expected to undergo amplification through various mechanisms: magnetic flux conservation (Woltjer, 1964), turbulent dynamo action (Duncan & Thompson, 1992; Thompson & Duncan, 1993; Masada et al., 2022), magnetorotational instability inside the PNS (Obergaulinger et al., 2009; Akiyama et al., 2003; Masada et al., 2015), and/or steady accretion shock instability of the post-bounce core-collapse supernova environment (Blondin et al., 2003). On the other hand, it is thought that the NS magnetic field decays over a relatively long time scale, $\gg 10~{}{\rm kyr}$, due to Ohmic diffusion, ambipolr diffusion, and Hall drift (e.g., Goldreich & Reisenegger, 1992; Takahashi et al., 2011; Skiathas & Gourgouliatos, 2024).
Supernova fallback has also been proposed as a contributing factor in the magnetic field evolution, particularly in the formation of the CCOs (see e.g., Muslimov & Page, 1995; Young & Chanmugam, 1995; Geppert et al., 1999). The fallback phase starts several seconds after the core bounce at the time when the neutrino-driven wind subsides (see e.g., Janka et al., 2022). If the fallback accretion rate is sufficiently high, the PNS’s magnetic field will be submerged beneath a crust newly formed by the fallback material. This case corresponds to the CCO formation (Torres-Forné et al., 2016; Shigeyama & Kashiyama, 2018). This scenario is referred to as the hidden magnetic field scenario. The hidden magnetic field is also believed to be the energy source of giant flares observed in soft gamma-ray repeaters (Thompson & Duncan, 2001).
The fallback accretion rate remains highly uncertain. Based on numerical simulations of the supernova explosion (e.g., Zhang et al., 2008; Dexter & Kasen, 2013; Moriya et al., 2018), the fallback accretion rate is approximated to be (Metzger et al., 2018; Barrère et al., 2022)
@@
where $t$ stands for time, and $M_{\rm fb}$ and $t_{\rm fb}$ are the typical fallback mass and fallback time scale, respectively. The dependence of $\dot{M}_{\rm fb}\propto t^{-5/3}$ for $t\gg t_{\rm fb}$ is expected for an accretion of marginally gravitationally bound matter (Michel, 1988; Chevalier, 1989). The typical fallback mass is estimated to be $M_{\rm fb}\sim 10^{-(4-1)}~{}{\rm M_{\odot}}$, depending on the progenitor structure (see e.g., Ugliano et al., 2012; Ertl et al., 2016a, b). The fallback time $t_{\rm fb}$ typically ranges from $1~{}{\rm s}$ to $10^{3}~{}{\rm s}$ (Janka et al., 2022), depending on the mean density of the progenitor star layer from which the fallback mass $M_{\rm fb}$ originates (Metzger et al., 2018). Therefore, $\dot{M}_{\rm fb}$ spans a wide range of $(10^{-7}-10^{-1})~{}{\rm M_{\odot}~{}s^{-1}}$.
@@
Figure 1: A schematic view of the supernova fallback. The white circle in the left figure denotes the PNS with a radius of $r_{\rm PNS}$. gray arrows stand for the velocity vector of the free-falling gas, and black wavy lines with arrows indicate the neutrinos propagating outward. We denote the magnetospheric and shock radii as $r_{\rm M}$ and $r_{\rm sh}$, respectively, in terms of the spherical radius $r$. The overall structure of the supernova fallback can be divided into free-fall region ($r>r_{\rm sh}$), post-shock region ($r_{\rm M}<r<r_{\rm sh}$), and the PNS magnetosphere ($r_{\rm PNS}<r<r_{\rm M}$). The shadow in the post-shock region represents the gas density and temperature. The right two panels are enlarged views around $r=r_{\rm M}$. A new crust forms above the PNS magnetosphere when neutrino cooling is efficient (panel a), whereas no crust forms when neutrino cooling is inefficient (panel b).
@@
As illustrated in the left panel of Figure 1, we consider that the fallback accretion onto a magnetized PNS consists of three regions (Torres-Forné et al., 2016): the free-fall region ($r>r_{\rm sh}$), the post-shock region ($r_{\rm M}<r<r_{\rm sh}$), and the PNS magnetosphere ($r_{\rm PNS}<r<r_{\rm M}$). Here, we denote the magnetospheric and shock radii by $r_{\rm M}$ and $r_{\rm sh}$, respectively, in terms of the spherical radius $r$. We write the PNS radius as $r_{\rm PNS}$. For $r>r_{\rm sh}$, the gas accretes toward the PNS at a free-fall velocity. For $r_{\rm M}<r<r_{\rm sh}$, a strong shock forms and the fallback material slowly settles to the surface of $r=r_{\rm M}$. For $r_{\rm PNS}<r<r_{\rm M}$, the PNS’s magnetic field is confined by the post-shock flows. When $\dot{M}_{\rm fb}$ is very high, neutrino emission occurs at $r\sim r_{\rm M}$. Such neutrino emission releases the gravitational energy liberated at $r\sim r_{\rm M}$. The two panels on the right are enlarged views around $r=r_{\rm M}$. These panels describe the cases where neutrino cooling is effective (panel a) and ineffective (panel b). If the neutrino cooling is effective enough for the new crust to form, then the PNS magnetosphere is submerged. Additionally, if the PNS rotation is significant, a bipolar jet powered by the magnetic field (Parfrey & Tchekhovskoy, 2017) or neutrinos (Qian & Woosley, 1996) will emerge around the rotation axis.
@@
Neutrino emissions affect the dynamics of the supernova fallback. Chevalier (1989) found steady solutions in which the gravitational energy released at the PNS surface is lost through neutrino cooling in the post-shock region. The steady solution is later confirmed by one-dimensional general relativistic Boltzmann neutrino-hydrodynamic simulations by Akaho et al. (2024). Since these studies considered the non-magnetzed PNS, it is still not understood how the neutrino emission affects the dynamics of the supernova fallback onto a magnetized PNS.
@@
In three dimensions, convection can influence the submergence of the PNS magnetic fields through convective diffusion. Bernal et al. (2013) presents a set of three-dimensional magnetohydrodynamic (MHD) simulations examining the interaction between fallback matter and an isolated magnetic arcade, suggesting that the post-shock region is marginally convectively stable (see also Bernal et al., 2010). Therefore, one- and two-dimensional models that neglect the effects of convection remain reasonable. Since this pioneering three-dimensional study explored only a limited parameter space, comprehensive surveys using one- or two-dimensional models are important for identifying the conditions under which the submergence of the magnetic field occurs.
@@
This paper is organized as follows. In Section 2, we briefly review the steady solution without magnetic fields in Chevalier (1989), which provides a key concept of the fallback process. We will present the numerical methods and models in Section 3 and show the numerical results in Section 4. Section 5 is devoted to the discussion of the diversity of young isolated NSs (Section 5.1) and neutrino scattering and absorption (Section 5.2). Finally, we give our conclusion in Section 6. Hereafter, the speed of light $c$ and the gravitational constant $G$ are normalized to $1$ unless otherwise specified. We represent the PNS mass as $M_{\rm PNS}$ and the gravitational radius as $r_{\rm g}=M_{\rm PNS}$.
@@ 2 Overview of Chevalier 1989
We briefly review the steady solution of the supernova fallback onto a non-magnetized PNS (Chevalier, 1989), which provides a guideline of our analysis. Since neutrino cooling is effective only in the vicinity of the PNS surface, the post-shock flow will be adiabatic over most of the accretion flow. We consider situations in which radiation pressure dominates gas pressure because of high accretion rates. To take into account the effect, we take an effective specific heat ratio of $\Gamma=4/3$ for the equation of state. Additionally, since local thermodynamic equilibrium will be achieved in the flows, we regard the radiation and gas temperatures as equal. Thus, we calculate the gas temperature as $T=(3p/a_{\rm rad})^{1/4}$, where $a_{\rm rad}$ is the radiation constant.
@@
where $\dot{q}$ is the cooling rate due to the neutrino emission. Chevalier (1989) adopted $H=r_{\rm PNS}^{2}p/({M_{\rm PNS}\rho})=(12/49)r_{\rm PNS}$ for the fiducial width of the layer in which neutrino cooling works effectively. From Equations (2) - (5), we get $r_{\rm sh}$ as a function of $\dot{M}_{\rm fb}$.
@@
The left panel of Figure 2 shows the cooling rate as a function of the gas temperature. We consider two processes relevant to the fallback accretion (Itoh et al., 1989; Qian & Woosley, 1996):
@@ 3 Numerical Method
We numerically solve the GRMHD equations while considering neutrino cooling in Schwarzschild polar coordinates $(t,r,\theta,\phi)$. These simulations are one-dimensional in spherical symmetry. We use the numerical code UWABAMI (e.g., Takahashi & Ohsuga, 2017). We represent space-time and space components as Greek and Latin suffixes, respectively.
@@ 3.1 Basic equations
The numerical methods used to solve the above equations are as follows. We utilize a harmonic mean proposed by van Leer (1977) for the reconstruction and adopt the Harten–Lax–van Leer (HLL; Harten et al., 1983) scheme to evaluate the numerical flux. We perform the time integration of the advection term using an explicit second-order Runge-Kutta scheme. On the other hand, we implicitly solve the source term responsible for the neutrino cooling $Q^{\mu}$ with the Newton-Raphson method. Since the gas velocity in the region where neutrino cooling is pronounced is much smaller than the speed of light, we take $Q^{i}=0$. This approach enables us to solve the source term using the Newton-Raphson method for a single variable, $p$, alone, rather than for four variables, $p$ and $u^{i}$, thereby reducing computational cost. Although we do not solve the energy and momentum transport due to neutrino scattering and absorption, these processes do not affect our conclusions within the density range treated in the present study (see Section 5.2).
@@
Unit $[\rm G]$ $[\rm g~{}cm^{-3}]$ $[\rm M_{\odot}~{}s^{-1}]$ $[\rm s]$
BnMm5R10 00 $10^{2}$ $10^{-5}$ $3.48\times 10^{-2}$
BnMm4R10 00 $10^{3}$ $10^{-4}$ $3.49\times 10^{-2}$
BnMm3R10 00 $10^{4}$ $10^{-3}$ $3.48\times 10^{-2}$
BnMm2R10 00 $10^{5}$ $10^{-2}$ $3.48\times 10^{-2}$
@@
BnMm4R20 00 $10^{3}$ $10^{-4}$ $3.49\times 10^{-2}$
BnMm3R20 00 $10^{4}$ $10^{-3}$ $3.49\times 10^{-2}$
BnMm2R20 00 $10^{5}$ $10^{-2}$ $3.48\times 10^{-2}$

### v2

@@ Abstract
Central compact objects (CCOs) are a subclass of neutron stars with a dipole magnetic field strength considerably weaker than those of radio pulsars and magnetars. One possible explanation for such weak magnetic fields in the CCOs is the hidden magnetic field scenario, in which supernova fallback submerges the magnetosphere of a proto-neutron star beneath a newly formed crust. However, the fallback mass and time scale required for this submergence process remain uncertain. We perform one-dimensional general relativistic magnetohydrodynamic simulations of the supernova fallback onto a magnetized proto-neutron star, while considering neutrino cooling. In our simulations, the infalling material compresses the magnetic field and drives a strong shock. The shock initially expands outward, but eventually stalls and recedes as neutrino cooling becomes significant. After the shock stalls, the gas density above the magnetosphere increases rapidly, potentially leading to the formation of a new crust. To understand the shock dynamics, we develop semi-analytic models that describe the resulting magnetospheric and shock radii when the shock stalls. By comparing the fallback time scale with the shock stalling time scale, corresponding to the waiting time for the new crust formation, we derive a necessary condition for the submergence of the proto-neutron star’s magnetic field. Our results will provide guidance for investigating the diversity of young isolated neutron stars through multidimensional simulations.
@@ 1 Introduction
Based on multi-wavelength observations, young isolated neutron stars (NSs) in our galaxy with ages of $\lesssim 1-10~{\rm kyr}$ are classified as magnetars, radio pulsars, and central compact objects (CCOs). The central engines powering these objects are believed to be the magnetic energy, rotation energy, and thermal energy of the NSs, respectively (e.g., see reviews by Enoto2019; Borghese2023). Their characteristic values of the dipole magnetic field strength at the NS surface are considerably different: $10^{14}~{\rm G}$ for magnetars, $10^{12-13}~{\rm G}$ for radio pulsars, and $10^{11}~{\rm G}$ for CCOs. However, the origin of such diversity remains unclear.
What determines the magnetic field strength of NSs is an open question (e.g., see reviews by Igoshev2021). When a proto-NS (PNS) forms in a core-collapse supernova, the magnetic field of the progenitor core is expected to undergo amplification through various mechanisms: magnetic flux conservation (Woltjer1964), turbulent dynamo action (Duncan1992; Thompson1993; Raynaud2020; Masada2022), magnetorotational instability inside the PNS (Akiyama2003; Obergaulinger2009; Masada2015; Reboul2021), and/or steady accretion shock instability of the post-bounce core-collapse supernova environment (Blondin2003; Endeve2010; Endeve2012). On the other hand, it is thought that the NS magnetic field decays over a relatively long time scale, $\gg 10~{\rm kyr}$, due to Ohmic diffusion, ambipolr diffusion, and Hall drift (e.g., Goldreich1992; Takahashi2011; Skiathas2024).
Supernova fallback has also been proposed as a contributing factor in the magnetic field evolution, particularly in the formation of the CCOs (Bernal2010; Ho2011; Shabaltas2012). The fallback phase starts several seconds or more after the core bounce at the time when the neutrino-driven wind subsides (see e.g., Thompson2001_wind; Janka2022; Wang2023; Wang2024). If the fallback accretion rate is sufficiently high, the PNS’s magnetic field will be submerged beneath a crust newly formed by the fallback material (Young1995; Muslimov1995; Geppert1999). This case corresponds to the CCO formation (Torres2016; Shigeyama2018). This scenario is referred to as the hidden magnetic field scenario. Depending on the fallback mass and the magnetic field strength, the hidden magnetic fields are expected to re-emerge through the crust on a time scale of $10^{1-2}~{\rm kyr}$ or longer Ho2011; Vigano2012; Igoshev2016; Gourgouliatos2020; Igoshev2021_a; Dehman2023; Fraija2025. The hidden magnetic field is also believed to be the energy source of giant flares observed in soft gamma-ray repeaters (Thompson2001).
The fallback accretion rate remains highly uncertain. Based on numerical simulations of the supernova explosion (e.g., Zhang2008; Dexter2013; Moriya2018), the fallback accretion rate is approximated to be (Metzger2018; Barrere2022)
@@
where $t$ stands for time, and $M_{\rm fb}$ and $t_{\rm fb}$ are the typical fallback mass and fallback time scale, respectively. The dependence of $\dot{M}_{\rm fb}\propto t^{-5/3}$ for $t\gg t_{\rm fb}$ is expected for an accretion of marginally gravitationally bound matter (Michel1988; Chevalier1989). The typical fallback mass is estimated to be $M_{\rm fb}\sim 10^{-(4-1)}~{\rm M_{\odot}}$, depending on the progenitor structure (see e.g., Ugliano2012; Ertl2016a; Ertl2016b). The fallback time $t_{\rm fb}$ typically ranges from $1~{\rm s}$ to $10^{3}~{\rm s}$ (Janka2022), depending on the mean density of the progenitor star layer from which the fallback mass originates (Metzger2018). Therefore, $\dot{M}_{\rm fb}$ spans a wide range of $(10^{-7}-10^{-1})~{\rm M_{\odot}~s^{-1}}$.
@@
Figure 1: A schematic view of the supernova fallback. The white circle in the left figure denotes the PNS with a radius of $r_{\rm PNS}$. The gray arrows stand for the velocity vector of the free-falling gas, and black wavy lines with arrows indicate the neutrinos propagating outward. We denote the magnetospheric and shock radii as $r_{\rm M}$ and $r_{\rm sh}$, respectively, in terms of the spherical radius $r$. The overall structure of the supernova fallback can be divided into free-fall region ($r>r_{\rm sh}$), post-shock region ($r_{\rm M}<r<r_{\rm sh}$), and the PNS magnetosphere ($r_{\rm PNS}<r<r_{\rm M}$). The shadow in the post-shock region represents the gas density and temperature. The right two panels are enlarged views around $r=r_{\rm M}$. A new crust forms above the PNS magnetosphere when neutrino cooling is efficient (panel a), whereas no crust forms when neutrino cooling is inefficient (panel b).
@@
As illustrated in the left panel of Figure 1, we consider that the fallback accretion onto a magnetized PNS consists of three regions (Torres2016): the free-fall region ($r>r_{\rm sh}$), the post-shock region ($r_{\rm M}<r<r_{\rm sh}$), and the PNS magnetosphere ($r_{\rm PNS}<r<r_{\rm M}$). Here, we denote the magnetospheric and shock radii by $r_{\rm M}$ and $r_{\rm sh}$, respectively, in terms of the spherical radius $r$. We write the PNS radius as $r_{\rm PNS}$. For $r>r_{\rm sh}$, the gas accretes toward the PNS at a free-fall velocity. For $r_{\rm M}<r<r_{\rm sh}$, a strong shock forms and the fallback material slowly settles to the surface of $r=r_{\rm M}$. For $r_{\rm PNS}<r<r_{\rm M}$, the PNS’s magnetic field is confined by the post-shock flows. When $\dot{M}_{\rm fb}$ is very high, neutrino emission occurs at $r\sim r_{\rm M}$. Such neutrino emission releases the gravitational energy liberated at $r\sim r_{\rm M}$. The two panels on the right are enlarged views around $r=r_{\rm M}$. These panels describe the cases where neutrino cooling is effective (panel a) and ineffective (panel b). If the neutrino cooling is effective enough for the new crust to form, then the PNS magnetosphere is submerged. Additionally, if the PNS rotation is significant, a bipolar jet powered by the magnetic field (Parfrey2017) or neutrinos (Qian1996) will emerge around the rotation axis.
@@
Neutrino emissions affect the dynamics of the supernova fallback. Chevalier1989 found steady solutions in which the gravitational energy released at the PNS surface is lost through neutrino cooling in the post-shock region (see also Houck1991). The steady solution is later confirmed by one-dimensional general relativistic Boltzmann neutrino-hydrodynamic simulations by Akaho2024. Since these studies considered the non-magnetzed PNS, it is still not understood how the neutrino emission affects the dynamics of the supernova fallback onto a magnetized PNS.
@@
In three dimensions, convection can influence the submergence of the PNS magnetic fields through convective diffusion. Bernal2013 presents a set of three-dimensional magnetohydrodynamic (MHD) simulations examining the interaction between fallback matter and an isolated magnetic arcade, suggesting that the post-shock region is marginally convectively stable (see also Bernal2010). Therefore, one- and two-dimensional models that neglect the effects of convection remain reasonable. Since this pioneering three-dimensional study explored only a limited parameter space, comprehensive surveys using one- or two-dimensional models are important for identifying the conditions under which the submergence of the magnetic field occurs.
@@
This paper is organized as follows. In Section 2, we briefly review the steady solution without magnetic fields in Chevalier1989, which provides a key concept of the fallback process. We will present the numerical methods and models in Section 3 and show the numerical results in Section 4. Section 5.1 is devoted to the discussion of the diversity of young isolated NSs. We discuss the limitations of our models in Section 5.2. Finally, we give our conclusion in Section 6. Hereafter, the speed of light $c$ and the gravitational constant $G$ are normalized to $1$ unless otherwise specified. We represent the PNS mass as $M_{\rm PNS}$ and the gravitational radius as $r_{\rm g}=M_{\rm PNS}$.
@@ 2 Overview of Chevalier 1989
We briefly review the steady solution of the supernova fallback onto a non-magnetized PNS (Chevalier1989), which provides a guideline of our analysis. Since neutrino cooling is effective only in the vicinity of the PNS surface, the post-shock flow will be adiabatic over most of the accretion flow. We consider situations in which radiation pressure dominates gas pressure because of high accretion rates. To take into account the effect, we take an effective specific heat ratio of $\Gamma=4/3$ for the equation of state (EoS). Additionally, since local thermodynamic equilibrium will be achieved in the flows, we regard the radiation and gas temperatures as equal. Thus, we calculate the gas temperature as $T=(3p/a_{\rm rad})^{1/4}$, where $a_{\rm rad}$ is the radiation constant. The electron-positron pair creation affects the pressure in reality, but we ignore it for simplicity. We will discuss the effect in Section 5.2.
@@
where $\dot{q}$ is the cooling rate due to the neutrino emission. Chevalier1989 adopted $H=r_{\rm PNS}^{2}p/({M_{\rm PNS}\rho})=(12/49)r_{\rm PNS}$ for the fiducial width of the layer in which neutrino cooling works effectively. From Equations (2) - (5), we get $r_{\rm sh}$ as a function of $\dot{M}_{\rm fb}$.
@@
The left panel of Figure 2 shows the cooling rate as a function of the gas temperature. We consider two processes relevant to the fallback accretion (Itoh1989; Qian1996):
@@ 3 Numerical Method
We numerically solve the GRMHD equations while considering neutrino cooling in Schwarzschild polar coordinates $(t,r,\theta,\phi)$. These simulations are one-dimensional in spherical symmetry, with the magnetic field depending only on the radius (Cumming2001). We use the numerical code UWABAMI (e.g., Takahashi2017). We represent space-time and space components as Greek and Latin suffixes, respectively.
@@ 3.1 Basic equations
The numerical methods used to solve the above equations are as follows. We utilize a harmonic mean proposed by vanLeer1977 for the reconstruction and adopt the Harten–Lax–van Leer (HLL; Harten1983) scheme to evaluate the numerical flux. We perform the time integration of the advection term using an explicit second-order Runge-Kutta scheme. On the other hand, we implicitly solve the source term responsible for the neutrino cooling $Q^{\mu}$ with the Newton-Raphson method. Since the gas velocity in the region where neutrino cooling is pronounced is much smaller than the speed of light, we take $Q^{i}=0$. This approach enables us to solve the source term using the Newton-Raphson method for a single variable, $p$, alone, rather than for four variables, $p$ and $u^{i}$, thereby reducing computational cost. Although we do not solve the energy and momentum transport due to neutrino scattering and absorption, these processes do not affect our conclusions within the density range treated in the present study (see Appendix A).
@@
Unit $\rm G] [$$\rm g~cm^{-3}] [$$\rm M_{\odot}~s^{-1}] [$$\rm s]$
BnMm5R10 0 $10^{2}$ $10^{-5}$ $3.48\times 10^{-2}$
BnMm4R10 0 $10^{3}$ $10^{-4}$ $3.49\times 10^{-2}$
BnMm3R10 0 $10^{4}$ $10^{-3}$ $3.48\times 10^{-2}$
BnMm2R10 0 $10^{5}$ $10^{-2}$ $3.48\times 10^{-2}$
@@
BnMm4R20 0 $10^{3}$ $10^{-4}$ $3.49\times 10^{-2}$
BnMm3R20 0 $10^{4}$ $10^{-3}$ $3.49\times 10^{-2}$
BnMm2R20 0 $10^{5}$ $10^{-2}$ $3.48\times 10^{-2}$